OpenStudy (anonymous):

. Find the area of the triangle. (Points : 2) 48.45 in2 20.4 in2 40.8 in2 81.6 in2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = (b*h)/2 A = (10.2*8)/2 A = ???

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

43.2

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

10.2 times 8 = ?

OpenStudy (anonymous):

oops 40.8

jimthompson5910 (jim_thompson5910):

you got it